A resting systolic blood pressure (BP) of 120 mm Hg is stated to be the population...

  1. A resting systolic blood pressure (BP) of 120 mm Hg is stated to be the population mean for women. This was based on research conducted by several researchers on women from the healthy general public. Consider 20 first-year resident female doctors drawn at random. The resting systolic BP measured using an electronic sphygmomanometer are listed below. Is the sample significantly different from the population?
    1. State the null and alternative hypotheses
    2. Set the criteria for a decision: two-tailed test with a = .05
    3. Compute the test statistic using SPSS
    4. What is the effect size (you will need to compute Cohen’s d from data given in the SPSS output table)

d= M- μ / SD  

  1. Write the Results section for your findings. Include the descriptive statistics, type of statistical test and results of the test, effect size, Cohen’s d, and the confidence interval. Use the example I’ve given you.

Resting systolic BP (each cell is an individual score):

128

127

111

132

118

115

149

122

144

142

139

119

133

140

136

129

132

131

126

128

Print the SPSS output table and write the answers to the above question on that page to be graded.
Answer #1

ans ) a)

b ) if the p value of the test statistic will be less than 0.05 then we reject the null hypothesis otherwise we accept the null hypothesis

c )

using spss the following is the output

One-Sample Statistics

N

Mean

Std. Deviation

Std. Error Mean

data

20

130.0500

9.96032

2.22719

One-Sample Test

Test Value = 0

t

df

Sig. (2-tailed)

Mean Difference

95% Confidence Interval of the Difference

Lower

Upper

data

58.392

19

.000

130.05000

125.3884

134.7116

the test statistic t = 58.392

p value = 0.000

d ) d= M- μ / SD   = ( 130.05 -120 ) /9.96 = 1.009

e )since the p value is less than 0.05 so we reject the null hypothesis and conclude that the population mean of women BP is not 120